# Near-optimal discrimination of displaced squeezed binary signals using displacement, inverse-squeezing, and photon-number-resolving detection

### 17 Mar 2026 v2

@@ Abstract
We propose an inverse-squeezing Kennedy receiver for discriminating binary phase-shift-keyed displaced squeezed vacuum states. The receiver combines a Kennedy-type nulling displacement, an orthogonally oriented inverse-squeezing operation and photon-number-resolving detection with a maximum-a-posteriori threshold rule. Its key mechanism is that the inverse-squeezing stage converts transmitter-side squeezing into enhanced photon-number contrast, or equivalently an effective coherent-state separation gain, that can be directly exploited at the measurement stage. Under ideal equal-prior conditions, the receiver surpasses the standard quantum limit for squeezed-state binary phase-shift keying at approximately $N\approx 0.3$, outperforms the Helstrom bound of coherent-state binary phase-shift keying at approximately $N\approx 0.4$, and reaches the 1% error level near $N\approx 0.6$. We further analyze its performance under realistic imperfections, including finite detector efficiency, dark counts, channel phase diffusion, receiver thermal noise and transmission loss. The results show that adaptive thresholding preserves robust performance against detector and noise imperfections over practical parameter ranges, whereas transmission loss progressively suppresses the squeezing-enabled advantage. These findings indicate that, for the fixed source parametrization adopted in this work, the proposed receiver is most advantageous in the low-loss regime, especially at low source energies.
preprint: APS/123-QED
I Introduction
@@
Motivated by this observation, the key open question is how to design a practical receiver that can efficiently convert the squeezing resource of the transmitted S-BPSK signals into an experimentally accessible discrimination advantage. In this work, we address this problem by proposing the inverse-squeezing Kennedy receiver (IS-Kennedy), which combines a displacement operation, an orthogonally oriented inverse-squeezing (IS) operation, and photon-number-resolving (PNR) detection with a maximum-a-posteriori (MAP) threshold rule. The basic idea is to map the incoming displaced squeezed-state alphabet into a more distinguishable effective coherent-state on–off-keying pair, thereby increasing the phase-space separation and reducing the overlap of the corresponding photon-number statistics. Under ideal conditions, we show that the error probability of the proposed receiver always remains within a factor of two of the S-BPSK Helstrom bound for equal priors, while outperforming coherent-state benchmarks in the low-energy regime. We further analyze its robustness under realistic imperfections, including finite detector efficiency, dark counts, phase diffusion, and receiver thermal noise.
The structure of this paper is as follows: In Sec. II, we review the signal model and benchmark limits for S-BPSK discrimination. In Sec. III, we introduce the IS-Kennedy receiver and analyze its operation and performance under ideal conditions. In Sec. IV, we study its performance in the presence of nonideal detectors and environmental noise. Finally, Sec. V concludes the paper.
II Channel model and performance benchmarks
@@
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Appendix A). And the resulting error probability is
@@
Figure 1: The ratios of $P_{\text{HB}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)$, $P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)$ and $P_{\text{SQL}}^{\text{C-BPSK}}\left(N\right)$ to $P_{\text{HB}}^{\text{C-BPSK}}\left(N\right)$ vs. input state energy N. Here, $P_{\text{SQL}}^{\text{C-BPSK}}\left(N\right)=\frac{1}{2}\left[1-\mathrm{erf}\left(\sqrt{2\cdot N}\right)\right]$ and $P_{\text{HB}}^{\text{C-BPSK}}\left(N\right)=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$.
@@
III inverse-squeezing Kennedy receiver under ideal conditions
@@
Figure 3: Phase-space representations (a–c) and Fock-basis populations (d–f) of the input states $\rho_{i}$ in the IS-Kennedy for mean photon number $N=1.0$ and squeezing fraction $\beta=\beta_{\text{opt}}$. Panels (a) and (d) show the input states $\rho_{i}$; panels (b) and (e) show the states after the nulling displacement $\sigma_{i}=D(\alpha)\rho_{i}D^{\dagger}(\alpha)$. Panels (c) and (f) show the states after inverse-squeezing $\zeta_{i}=U\rho_{i}U^{\dagger}$ with $U=S(-r)D(\alpha)$.
@@
where $P_{\text{err}}^{\text{K},\text{ideal}}$ is the error probability of the conventional Kennedy receiver under ideal conditions [22]. This implies that, under matched squeezing, the IS-Kennedy is equivalent to mapping an S-BPSK signal with mean photon number $N$ to a C-BPSK signal with mean photon number $N_{\text{eff}}$ via the IS module, followed by Kennedy nulling and on-off detection.
@@
Figure 4: Performance of the IS-Kennedy under ideal conditions. (a) Error probability of the IS-Kennedy. (b) Ratio (in dB) of the benchmarks $P_{\text{HB}}^{\text{C-BPSK}}$, $P_{\text{SQL}}^{\text{C-BPSK}}$, $P_{\text{HB}}^{\text{S-BPSK}}$, $P_{\text{SQL}}^{\text{S-BPSK}}$, and $P_{\text{err}}^{\text{K},\text{ideal}}$ relative to $P_{\text{err}}^{\text{ISK},\text{ideal}}$.
@@
IV The performance of IS-Kennedy under imperfect conditions
IV.1 Imperfect PNR detector with detection efficiency, dark count, and finite resolution
@@
Considering an imperfect detection efficiency $\eta\in(0,1]$ and a mean dark count rate $\nu\geq 0$, the POVM operators are described by [3, 19]:
@@
$P(n|i)=\text{Tr}\left\{\zeta_{i}\Pi_{n}^{\left(\eta,\nu,\text{M}\right)}\right\}=\left\{\begin{aligned} &\text{Poiss}\left(n;\eta\mu_{i}+\nu\right),\quad n<\text{M}-1,\\ &\sum_{m=M}^{\infty}\text{Poiss}\left(m;\eta\mu_{i}+\nu\right),\quad n=\text{M},\end{aligned}\right$. (28)
@@
Figure 5: Performance analysis of the IS-Kennedy receiver under imperfect detection conditions. (a) Error probability of IS-Kennedy with SPD for varying detection efficiencies $\eta\in\left\{1.0,0.8,0.4\right\}$ and extremely small dark counts ($\nu=10^{-9}$). (b) Error probability of IS-Kennedy with PNR(2) under different dark count rates $\nu$ (with $\eta=1.0$). The “step-like” behavior in the error curves arises from the discrete nature of the optimal decision threshold. (c) The optimal detection threshold $n_{\text{th}}^{*}$ as a function of signal energy for different dark count rates, calculated via Eq. (30). The jumps in threshold correspond exactly to the inflection points observed in panels (b) and (d). (d) The performance ratio (in dB) of the IS-Kennedy with PNR(10) relative to the S-BPSK SQL. The oscillatory behavior reflects the discrete adjustment of $n_{\text{th}}^{*}$ to balance signal detection against dark count noise.
@@
Given that the dark count rate of practical PNR detectors is typically much less than unity ($\nu\ll 1$[13], the dominant contribution to the summation comes from the first term, $n=\text{M}$. Thus, the saturation error probability can be approximated as:
@@
IV.2 Phase Diffusion Noise in communication channel
Phase diffusion noise is a common communication channel noise that increases the overlap between states, making them more difficult to distinguish [26, 10]. A schematic of a phase diffusion noise channel is shown in Fig. 6.
Figure 6: Schematic diagram of phase diffusion noise channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{pd}$ (mixed state) in the phase diffusion communication channel.
Figure 7: Phase-space representations of S-BPSK signal (mean photon number $N=1.0$ and phase diffusion strength $\sigma=0.5$). (a) S-BPSK $\rho_{i}$ without phase diffusion noise; (b) states $\rho_{i}^{\text{pd}}$ with phase diffusion noise; (c) states $\zeta_{i}$ after IS-Kennedy processing.
@@ IV.2 Phase Diffusion Noise in communication channel
$\left|\psi_{i,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{i,\phi}\right\rangle=S\left(-r\right)D\left((-1)^{i}\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (36)
@@
$P(n|i)=\text{Tr}\left\{\Pi_{n}^{\text{ideal}}\cdot\zeta_{i}\right\}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot\text{Tr}\left\{\Pi_{n}^{\text{ideal}}\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\}$ (37)
Figure 8: Simulation results of IS-Kennedy under phase diffusion noise ($\sigma=0.1$): (a) Error probability, (b) Ratio (in dB) with a benchmark of $P_{\text{SQL}}^{\text{S-BPSK}}$ (Eq. (65)), (c) Optimal threshold $n_{th}^{*}$ and (d) Fock-basis population of measured states $\zeta_{i}=U\rho_{i}^{\text{pd}}U^{\dagger}$ by PNR(M) for $N=1.5$. Note: the SQL and HB of S-BPSK under phase diffusion condition is rewritten as Eq. (65) and Eq. (67).
By substituting Eq. (37) into the MAP rule in Eq. (16), we numerically determine the optimal decision rule and the corresponding error probability under phase diffusion noise. We do not have a general closed-form proof that the likelihood ratio remains monotonic in $n$ for arbitrary $\sigma$; therefore, the threshold reported below should be understood as a numerically verified MAP-equivalent threshold over the explored parameter range.
Fig. 8(c) shows that the optimal threshold $n_{\text{th}}^{*}$ under phase diffusion ($\sigma=0.1$) exhibits a staircase-like dependence on the mean photon number $N$. In the absence of PNR resolution limits, $n_{\text{th}}^{*}$ increases predominantly in unit steps, while some even thresholds remain optimal only over very narrow intervals of $N$, leading to unequal step widths compared with the dark-count case in Sec. IV.A. This behavior originates from the non-Gaussianity induced by phase diffusion: after the IS-Kennedy processing, $\zeta_{i}$ cannot be perfectly nulled to (near-)vacuum and its photon-number distribution exhibits noticeable even-odd modulation (Fig. 8(d)). As $N$ varies, this modulation causes the likelihood function $L\left(n\right)$ to shift non-uniformly across adjacent photon numbers, so that certain even thresholds can be optimal only over short $N$ intervals. Consequently, $n_{\text{th}}^{*}$ still changes mainly by +1, but with uneven plateau lengths. We verified this by refining the sampling of $N$, which reveals short-lived plateaus at some even $n_{\text{th}}^{*}$ values that may be missed under coarser sampling.
@@
IV.3 Thermal noise in the receiver
@@
Figure 9: Schematic of the IS-Kennedy under receiver thermal noise. The processed states $\zeta_{i}$ impinging on the PNR(M) are contaminated by receiver thermal noise and become $\zeta_{i}^{\prime}$.
Consistent with Ref. [41, 40], we assume that the thermal-noise contamination occurs immediately before the PNR detector, as illustrated in Fig. 9.
Note (coherent state $\Rightarrow$ displaced thermal state): Under the channel $\Phi_{n_{t}}$ in Eq. (38), a coherent-state input becomes a displaced thermal state. Specifically, using the Weyl relation $D(\lambda)D(\alpha)=e^{(\lambda\alpha^{\ast}-\lambda^{\ast}\alpha)/2}D(\lambda+\alpha)$, one verifies the displacement covariance
$\Phi_{n_{t}}\!\big(D(\alpha)\rho D^{\dagger}(\alpha)\big)=D(\alpha)\Phi_{n_{t}}(\rho)D^{\dagger}(\alpha).$ (40)
Moreover, by definition of Eq. (38), $\Phi_{n_{t}}(|0\rangle\langle 0|)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,|\lambda\rangle\langle\lambda|$ is exactly a thermal state with mean photon number $n_{t}$, denoted as $\rho_{\mathrm{th}}(n_{t})$. Therefore, for $|\alpha\rangle=D(\alpha)|0\rangle$,
$\Phi_{n_{t}}\!\big(|\alpha\rangle\langle\alpha|\big)=\Phi_{n_{t}}\!\big(D(\alpha)|0\rangle\langle 0|D^{\dagger}(\alpha)\big)=D(\alpha)\rho_{\mathrm{th}}(n_{t})D^{\dagger}(\alpha).$ (41)
@@
Equivalently, by Eq. (41), one may write $\zeta_{0}^{\prime}=\rho_{\mathrm{th}}(n_{t})$ and $\zeta_{1}^{\prime}=D(2\gamma)\rho_{\mathrm{th}}(n_{t})D^{\dagger}(2\gamma)$.
@@
$P\left(n|i\right)=\text{Tr}\!\left(\Pi_{n}^{\text{ideal}}\,\zeta_{i}^{\prime}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\cdot f_{n_{t}}(\lambda)\cdot\text{Poiss}\left\{n;\mu_{i}(\lambda)\right\}$ (43)
where $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\gamma|^{2}$. Substituting Eq. (43) into the MAP rule in Eq. (16), we numerically obtain the optimal decision rule and the corresponding error probability in the presence of receiver thermal noise. A general closed-form monotonicity proof for the likelihood ratio is not available in this mixed-state setting; accordingly, the threshold values reported below are numerically verified MAP-equivalent thresholds within the explored parameter range.
Figure 10: Simulation results of IS-Kennedy with PNR(M) under thermal noise condition: (a) Error probability ($n_{t}=0.001$), (b) Error probability ($M=10$), (c) Ratio (in dB) of the benchmark $P_{\text{SQL}}^{\text{S-BPSK}}$ ($M=10$), (d) Optimal threshold ($M=10$), (e) Fock-basis population of $\zeta_{i}^{\prime}$ (the input energy $N=1.0$ and $n_{t}=0.001$), and (f) Maximum tolerable thermal noise $n_{t}^{\text{max}}$ for different photon number resolution M.
Fig. 10(a) and (b) depict the error probability of the IS-Kennedy as a function of the mean signal photon number, under different PNR detector resolutions and various thermal noise strengths. The error probability curves exhibit an oscillatory pattern and decrease overall with increasing signal energy, which is a consequence of the continuous adjustment of the optimal threshold $n_{\text{th}}^{*}$ (see panel (d)). This trend continues until $n_{\text{th}}^{*}$ reaches the maximum resolution M of the PNR detector, after which the error probability saturates and no longer varies with the mean photon number. As analyzed in Sec. IV.A, the saturated error probability is dominated by the false alarm probability $P_{\text{FA}}$ (Eq. (20)), and when the signal energy tends to infinity, the optimal threshold must be equal to M. Therefore, the saturated error probability is:
@@
Fig. 10(c) shows the ratio (in dB) of the IS-Kennedy error probability to the SQL of standard BPSK versus the mean signal photon number under different noise strengths. It can be observed that the IS-Kennedy can still surpass the SQL at certain noise levels. For instance, when $n_{t}=10^{-3}$, the error probability of the IS-Kennedy with a PNR(M) detector is lower than the SQL by -6.34dB (M=1), -13.43dB (M=2), -20.79dB (M=3), and -28.27dB (M=4).
@@
IV.4 Transmission loss in the communication channel
A practically important limitation of squeezed-state communication is transmission loss. In contrast to the ideal analysis of Sec. III, propagation through a lossy channel does not merely reduce the signal displacement, but also degrades the squeezing resource itself by mixing the signal mode with an environmental vacuum mode. Equivalently, in the Stinespring representation, a pure-loss bosonic channel can be modeled by a beam splitter of transmittance $T$ that couples the signal to an environmental vacuum state [10, 37].
@@ IV.4 Transmission loss in the communication channel
Figure 11: Schematic diagram of pure-loss channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{\text{loss}}$ (mixed state) in the pure-loss communication channel, which $T$ is the equivalent transmittance.
1. Input Gaussian-state description.
@@
2. Pure-loss channel as a Gaussian map.
We model propagation by a pure-loss bosonic channel with transmittance $T\in[0,1]$ (see Fig. 11). At the operator level, the channel can be represented as
@@
These expressions have a clear physical meaning. When $T\to 1$, one has $r_{T}\to r$ and $n_{T}\to 0$, so the ideal case is recovered. When $0<T<1$, the effective squeezing is reduced from $r$ to $r_{T}$, while a nonzero thermal component $n_{T}$ appears. When $T\to 0$, the signal is completely extinguished and both hypotheses collapse to vacuum. Hence, transmission loss affects the receiver in two distinct ways: it attenuates the displacement and, at the same time, destroys the purity of the squeezing resource.
4. Matched IS-Kennedy processing after a lossy channel.
In the ideal receiver of Sec. III, the nulling displacement $D(\alpha)$ and the inverse-squeezing S(r)S(-r) are matched to the transmitted state. After a lossy channel, the received hypotheses are centered at $\pm\sqrt{T}\alpha$, rather than at $\pm\alpha$, and exhibit residual squeezing $r_{T}$ rather than $r$. Therefore, the matched version of the IS-Kennedy in the lossy scenario should use
@@
the two output hypotheses impinging on the PNR detector become (an OOK pair composed of a thermal state and a displaced thermal state)
@@
$P(n|i=0)=\int_{\mathbb{C}}\text{d}^{2}\lambda\cdot f_{n_{T}}(\lambda)\cdot\mathrm{Poiss}(n;|\lambda|^{2}),$ (56)
$P(n|i=1)=\int_{\mathbb{C}}\text{d}^{2}\lambda\cdot f_{n_{T}}(\lambda)\cdot\mathrm{Poiss}(n;|\lambda+2\gamma_{T}|^{2}).$
These expressions are identical in structure to the thermal-noise formulas already used in Sec. IV.C, with the substitutions:
@@
Since the lossy case reduces to the same thermal-state decision structure with the substitutions in Eq. (57), the MAP decision rule and the corresponding error probability are evaluated numerically. As in Secs. IV.B and IV.C, we do not claim a general closed-form proof that the likelihood ratio is monotonic in $n$; the thresholds shown below are therefore numerically verified MAP-equivalent thresholds over the explored parameter range.
@@
The role of the PNR resolution is illustrated in Fig. 12(d). As analyzed in Sec. IV.C, when the signal energy increases, the optimal threshold increases to M, the miss detection probability $P_{\text{Mi}}$ decreases approaching 0, and the receiver’s error probability is dominated by false alarm probability $P_{\text{FA}}$, so that at very high input energy:
@@
Figure 13: Maximum link length $L_{\text{max}}$ of the benchmark $P_{\text{err}}^{\text{K,loss}}$ for different mean photon number $N$ and resolution M

### 12 May 2026 v3

@@ Abstract
We propose an inverse-squeezing Kennedy receiver for discriminating binary phase-shift-keyed displaced squeezed vacuum states. The receiver combines a Kennedy-type nulling displacement, an orthogonally oriented inverse-squeezing operation and photon-number-resolving detection with a maximum-a-posteriori threshold rule. Its key mechanism is that the inverse-squeezing stage converts transmitter-side squeezing into enhanced photon-number contrast, or equivalently an effective coherent-state energy gain, that can be directly exploited at the measurement stage. Under ideal equal-prior conditions, the receiver surpasses the standard quantum limit for squeezed-state binary phase-shift keying at approximately $N\approx 0.3$, outperforms the Helstrom bound of coherent-state binary phase-shift keying at approximately $N\approx 0.4$, and reaches the 1% error level near $N\approx 0.6$. We further analyze its performance under realistic imperfections, including finite detector efficiency, dark counts, channel phase diffusion, receiver thermal noise and transmission loss. The results show that adaptive thresholding preserves robust performance against detector and noise imperfections over practical parameter ranges, whereas transmission loss progressively suppresses the squeezing-enabled advantage. These findings indicate that, for the fixed source parametrization adopted in this work, the proposed receiver is most advantageous in the low-loss regime, especially at low source energies.
Introduction
@@
Motivated by this observation, the key open question is how to design a practical receiver that can efficiently convert the squeezing resource of the transmitted S-BPSK signals into an experimentally accessible discrimination advantage. In this work, we address this problem by proposing the inverse-squeezing Kennedy receiver (IS-Kennedy), which combines a displacement operation, an orthogonally oriented inverse-squeezing (IS) operation, and photon-number-resolving (PNR) detection with a maximum-a-posteriori (MAP) threshold rule. The basic idea is to map the incoming displaced squeezed-state alphabet into a more distinguishable effective coherent-state on–off-keying pair, thereby increasing the phase-space separation and reducing the overlap of the corresponding photon-number statistics. Under ideal conditions, we show that the error probability of the proposed receiver always remains within a factor of two of the S-BPSK Helstrom bound for equal priors, while outperforming coherent-state benchmarks in the low-energy regime. We further analyze its robustness under realistic imperfections, including finite detector efficiency, dark counts, phase diffusion, receiver thermal noise, and transmission loss.
Results
Channel model and performance benchmarks
@@ Results
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Supplementary Information I). And the resulting error probability is
@@
Inverse-squeezing Kennedy receiver under ideal conditions
@@
where $P_{\text{err}}^{\text{K},\text{ideal}}$ is the error probability of the conventional Kennedy receiver under ideal conditions [5]. This implies that, under matched squeezing, the IS-Kennedy is equivalent to mapping an S-BPSK signal with mean photon number $N$ to a C-BPSK signal with mean photon number $N_{\text{eff}}$ via the IS module, followed by Kennedy nulling and on-off detection. That is, the IS-Kennedy converts the squeezing resources of S-BPSK signals into higher coherent-state energy advantage ($N\to N_{\text{eff}}=N(N+1)$) through the IS module, thereby achieving a lower error probability. Compared to quantum receivers for coherent state discrimination, the energy gain brought by the IS-Kennedy is
$G=10\cdot\log_{10}\left(\frac{N_{\text{eff}}}{N}\right)=10\cdot\log_{10}\left(N+1\right)\quad(\text{dB})$ (24)
@@
Imperfect PNR detector with detection efficiency, dark count, and finite resolution
@@
Considering an imperfect detection efficiency $\eta\in(0,1]$ and dark counts $\nu\geq 0$, the POVM operators are described by [32, 33]:
@@
$P(n|i)=\text{Tr}\left\{\zeta_{i}\Pi_{n}^{\left(\eta,\nu,\text{M}\right)}\right\}=\left\{\begin{aligned} &\text{Poiss}\left(n;\eta\mu_{i}+\nu\right),\quad n\leq\text{M}-1,\\ &\sum_{m=M}^{\infty}\text{Poiss}\left(m;\eta\mu_{i}+\nu\right),\quad n=\text{M},\end{aligned}\right$. (29)
@@
Given that the dark count rate of practical PNR detectors is typically much less than unity ($\nu\ll 1$) [34], the dominant contribution to the summation comes from the first term, $n=\text{M}$. Thus, the saturation error probability can be approximated as:
@@
Phase Diffusion Noise in communication channel
Phase diffusion noise is a common communication channel noise that increases the overlap between states, making them more difficult to distinguish [37, 29, 38]. A schematic of a phase diffusion noise channel is shown in Fig. 6.
@@
$\left|\psi_{i,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{i,\phi}\right\rangle=S\left(-r\right)D\left((-1)^{i+1}\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (37)
@@
$P(n|i)=\text{Tr}\left\{\Pi_{n}^{(\text{M})}\cdot\zeta_{i}\right\}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot\text{Tr}\left\{\Pi_{n}^{(\text{M})}\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\},$ (38)
where
$\Pi_{n}^{\left(\text{M}\right)}=\left\{\begin{aligned} &\Pi_{n}^{\text{ideal}},\quad n=0,1,\cdots,\text{M}-1,\\ &\sum_{m=M}^{\infty}\Pi_{m}^{\text{ideal}},\quad n=\text{M}.\end{aligned}\right$. (39)
By substituting Eq. (38) into the MAP rule in Eq. (16), we numerically determine the optimal decision rule and the corresponding error probability under phase diffusion noise. The optimal threshold is obtained numerically as described in Sec. Methods.
Fig. 8(c) shows that the optimal threshold $n_{\text{th}}^{*}$ under phase diffusion ($\sigma=0.1$) exhibits a staircase-like dependence on the mean photon number $N$. In the absence of PNR resolution limits, $n_{\text{th}}^{*}$ increases predominantly in unit steps, while some even thresholds remain optimal only over very narrow intervals of $N$, leading to unequal step widths compared with the dark-count case in Sec. Results: Imperfect PNR detector with detection efficiency, dark count, and finite resolution. This behavior originates from the non-Gaussianity induced by phase diffusion: after the IS-Kennedy processing, $\zeta_{i}$ cannot be perfectly nulled to (near-)vacuum and its photon-number distribution exhibits noticeable even-odd modulation (Fig. 8(d)). As $N$ varies, this modulation causes the likelihood function $L\left(n\right)$ to shift non-uniformly across adjacent photon numbers, so that certain even thresholds can be optimal only over short $N$ intervals. Consequently, $n_{\text{th}}^{*}$ still changes mainly by +1, but with uneven plateau lengths. We verified this by refining the sampling of $N$, which reveals short-lived plateaus at some even $n_{\text{th}}^{*}$ values that may be missed under coarser sampling.
@@
Thermal noise in the receiver
@@
Consistent with Ref. [40, 41], we assume that the thermal-noise contamination occurs immediately before the PNR detector, as illustrated in Fig. 9. Therefore the HB and SQL of received states $\rho_{i}$ remain unchanged.
@@
Because the thermal-noise channel $\Phi_{n_{t}}$ in Eq. (41) is displacement covariant (namely, applying a displacement before the channel is equivalent to applying the same displacement after the channel), Eq. (42) may write
$\zeta_{0}^{\prime}=\rho_{\mathrm{th}}(n_{t}),\quad\zeta_{1}^{\prime}=D(2\gamma)\rho_{\mathrm{th}}(n_{t})D^{\dagger}(2\gamma).$ (43)
@@
$P\left(n|i\right)=\text{Tr}\!\left(\Pi_{n}^{\left(\text{M}\right)}\,\zeta_{i}^{\prime}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\cdot f_{n_{t}}(\lambda)\cdot\text{Poiss}\left\{n;\mu_{i}(\lambda)\right\}$ (44)
where $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\gamma|^{2}$. Substituting Eq. (44) into the MAP rule in Eq. (16), we numerically obtain the optimal decision rule and the corresponding error probability in the presence of receiver thermal noise. The optimal threshold is obtained numerically as described in Sec. Methods.
Fig. 10(a) and (b) depict the error probability of the IS-Kennedy as a function of the mean signal photon number, under different PNR detector resolutions and various thermal noise strengths. The error probability curves exhibit an oscillatory pattern and decrease overall with increasing signal energy, which is a consequence of the continuous adjustment of the optimal threshold $n_{\text{th}}^{*}$ (see panel (d)). This trend continues until $n_{\text{th}}^{*}$ reaches the maximum resolution M of the PNR detector, after which the error probability saturates and no longer varies with the mean photon number. As analyzed in Sec. Results: Imperfect PNR detector with detection efficiency, dark count, and finite resolution, the saturated error probability is dominated by the false alarm probability $P_{\text{FA}}$ (Eq. (20)), and when the signal energy tends to infinity, the optimal threshold must be equal to M. Therefore, the saturated error probability is:
@@
Fig. 10(c) shows the ratio (in dB) of the IS-Kennedy error probability to the SQL of S-BPSK versus the mean signal photon number under different noise strengths. It can be observed that the IS-Kennedy can still surpass the SQL at certain noise levels. For instance, when $n_{t}=10^{-3}$, the error probability of the IS-Kennedy with a PNR(M) detector is lower than the SQL by -6.34dB (M=1), -13.43dB (M=2), -20.79dB (M=3), and -28.27dB (M=4).
@@
Transmission loss in the communication channel
A practically important limitation of squeezed-state communication is transmission loss. In contrast to the ideal analysis of Sec. Results: inverse-squeezing Kennedy receiver under ideal conditions, propagation through a lossy channel does not merely reduce the signal displacement, but also degrades the squeezing resource itself by mixing the signal mode with an environmental vacuum mode. Equivalently, in the Stinespring representation, a pure-loss bosonic channel can be modeled by a beam splitter of transmittance $T$ that couples the signal to an environmental vacuum state [29, 4].
@@
We model propagation by a pure-loss bosonic channel with transmittance $T\in[0,1]$ (see Fig. 11) [4]. At the operator level, the channel can be represented as
@@
These expressions have a clear physical meaning. When $T\to 1$, one has $r_{T}\to r$ and $n_{T}\to 0$, so the ideal case is recovered. When $0<T<1$, the effective squeezing is reduced from $r$ to $r_{T}$, while a nonzero thermal component $n_{T}$ appears. When $T\to 0$, the signal is completely extinguished and both states collapse to vacuum. Hence, transmission loss affects the receiver in two distinct ways: it attenuates the displacement and, at the same time, destroys the purity of the squeezing resource.
Under ideal conditions, the nulling displacement $D(\alpha)$ and the inverse-squeezing S(r)S(-r) are matched to the transmitted state. After a lossy channel, the received states are centered at $\pm\sqrt{T}\alpha$, rather than at $\pm\alpha$, and exhibit residual squeezing $r_{T}$ rather than $r$. Therefore, the matched version of the IS-Kennedy in the lossy scenario should use
@@
the two output states impinging on the PNR detector become (an OOK pair composed of a thermal state and a displaced thermal state)
@@
$P\left(n|i\right)=\text{Tr}\!\left(\Pi_{n}^{\left(\text{M}\right)}\,\zeta_{i}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\cdot f_{n_{T}}(\lambda)\cdot\text{Poiss}\left\{n;\mu_{i}(\lambda)\right\},$ (57)
where $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\gamma_{T}|^{2}$. These expressions are identical in structure to the thermal-noise formulas already used in Sec. Results: Thermal noise in the receiver, with the substitutions:
@@
Since the lossy case reduces to the same thermal-state decision structure with the substitutions in Eq. (58), the MAP decision rule and the corresponding error probability are evaluated numerically. The optimal threshold is obtained numerically as described in Sec. Methods.
@@
The role of the PNR resolution is illustrated in Fig. 12(d). As analyzed in Sec. Results: Thermal noise in the receiver, when the signal energy increases, the optimal threshold increases to M, the miss detection probability $P_{\text{Mi}}$ decreases approaching 0, and the receiver’s error probability is dominated by false alarm probability $P_{\text{FA}}$, so that at very high input energy: